OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -4) with a slope of 1/2.

OpenStudy (anonymous):

\[y+4=\frac{1}{2}(x-2)\] is one form

OpenStudy (karatechopper):

where r the choices?

OpenStudy (anonymous):

\[y=\frac{1}{2}x-5\] could be another.

OpenStudy (aroub):

y=mx+b , m=1/2 , ( 2,-4) -4=1/2(2)+b -4=1+b -5=b y=1/2x-5